OpenStudy (anonymous):

PLEASE HELP ME Find the first error that Tony made in his work and circle it. (there may be more than one mistake!) -6 ÷ 2/3 = -6 * 2/3 = -6/1 * 2/3 = 12/3 = 4 would it be -6 ÷ 2/3 = -6 * 3/2 = -9?

OpenStudy (anonymous):

The mistake is in the beginning -6 / (2/3) = -6(2/3) does not equal each other Does that make sense?

OpenStudy (anonymous):

okay yes, that makes sense. So..i would have to divide -6 2/3?

OpenStudy (anonymous):

-6 / (2/3)?

OpenStudy (anonymous):

Instead of diving, you would want to multiply it. That goes along with the rest of the sequence -6 / 2/3 = -18/2 -6(2/3) = -12/3 They do not equal each other so that portion of the problem is incorrect.

OpenStudy (anonymous):

Also Toby made another mistake. It shouldn't be 12/3 = 4 It should be -12/3 = -4

OpenStudy (anonymous):

okay...do you mind typing it out like in order. like how the original problem is?

OpenStudy (anonymous):

The original problem from -6/(2/3) ? ? ?

OpenStudy (anonymous):

-6/(2/3) = -6(3)/2 = -18/2 = -9

OpenStudy (anonymous):

okay thank you! and thats the correct way he should have done it?

OpenStudy (anonymous):

From -6/(2/3) yeah :)

OpenStudy (anonymous):

okay cool thank you soo much! :D